OpenStudy (anonymous):

Write and equation of the line that passes through the given point and is perpendicular to the graph of the given equation: (5,7); y= 1/3x + 2

OpenStudy (inkyvoyd):

Okay. Do you have any idea where to start?

OpenStudy (anonymous):

Hint : if you have two straight lines and they are perpendicular Then M1 x M2 =-1 when M = Slope .

OpenStudy (inkyvoyd):

Here's the hint: The slope of a line perpendicular to another is the negative reciprocal. If the slope is m, then the slope of the perpendicular line is -1/m. Does that make sense?

OpenStudy (anonymous):

@inkyvoyd It sort of does... I understand that part but I don't get what comes next

OpenStudy (inkyvoyd):

Well, you are given a line. You want to know the slope of a line perpendicular to it. With the line you are given, what is the slope? Whatever the slope is, the slope of a line perpendicular to it is -1/m where m is the slope.

OpenStudy (anonymous):

@inkyvoyd The slope is already given so I did the reciprocol. So instead of the basic problem of (5,7); y= 1/3x + 2 ... So far I have 3x (+ or - somthing) I need to figure out the something

OpenStudy (inkyvoyd):

Wee, you need to be careful to take the negative. You have a slope m=-3. and you have to condition that (7)=(-3)(5)+b (I just substituted y, m, and x into y=mx+b, the equation of the slope-intercept line)

OpenStudy (anonymous):

@inkyvoyd SO my final answer would be y= -3x + 9 ?

OpenStudy (inkyvoyd):

No. How did you get that answer?

OpenStudy (anonymous):

@inkyvoyd Ack I knew it I'm a dumbarse I got it because -3 is the slope and using y= mx+b And -3 times 3 (5 -2) is -9

OpenStudy (inkyvoyd):

uhh, I'll work it out step by step y=mx+b m=-3 (x,y)=(5,7) (7)=-3(5)+b 7+15=b b=22

OpenStudy (anonymous):

@inkyvoyd Thanks, hopefully I can continue my homework correctly aha (:

OpenStudy (inkyvoyd):

okay - tell me your final answer :)

OpenStudy (anonymous):

@inkyvoyd y= -3x +22 ?? *prays that shes right*